# CADmium: Fine-Tuning Code Language Models for Text-Driven Sequential CAD Design

### v2

@@ Appendix D Qualitative Analysis of New Metrics
Appendix E Comparitive Qualitative Analysis
@@ Appendix F Licenses for External Assets
The DeepCAD dataset [Wu et al., 2021] is available under the MIT License, as indicated in its official repository. The ABC dataset [Koch et al., 2019], which is a source for DeepCAD, is also distributed under the MIT License. The Fusion 360 Reconstruction dataset [Willis et al., 2021], originating from the Autodesk Fusion 360 Gallery, is referenced in academic literature as being available for research purposes under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.
The language models leveraged include: Qwen2.5-Coder [Yang et al., 2024], open-source versions of which are provided under the Apache License 2.0. The GPT-4.1 model [OpenAI, 2025] is a proprietary offering from OpenAI, and its usage is governed by OpenAI’s prevailing Terms of Use, Usage Policies, and Sharing & Publication Policy. Gemma models [Team et al., 2025], developed by Google, are released under a custom Gemma license.
Key software utilized includes Blender [Blender Online Community, 2025], which is distributed under the GNU General Public License (GPL) version 3 or later. The Python library trimesh [Dawson-Haggerty et al., ], used for geometric computations, is available under the MIT License.

### v3

@@ Appendix D Qualitative Analysis of New Metrics
Appendix E Design Choices
Annotation Pipeline
Prior works like Text2CAD utilized ensembles of separate vision and language models. We consciously chose a single, state-of-the-art multimodal model (GPT-4.1) to unify the reasoning process. We hypothesized that a single model with strong instruction-following capabilities would reduce error propagation between the visual perception and text generation stages. We prioritized minimal JSON inputs over verbose serialized formats to maximize the token context window available for reasoning. Consistent with the methodology proposed in Text2CAD, we utilize 10 multi-view images as input for the annotation model.
Annotation Evaluation
For evaluating annotation quality, we employed Gemma-3 12B and Mistral Small 3.2 24B. We avoided using GPT-4 for evaluation to prevent self-preference bias, as our annotations were generated by GPT-4.1. By using distinct model families for evaluation, we aimed to decouple the generator’s bias from the evaluator’s preference.
Fine-tuning
We selected Qwen-2.5 Coder due to its demonstrated superiority in code generation benchmarks compared to general-purpose LLMs of similar size. For training, we utilized Low-Rank Adaptation (LoRA) with a rank of 64. This rank was chosen empirically; preliminary experiments suggested that lower ranks ($r=8,16$) were insufficient for the model to learn the strict syntax of the CAD JSON schema, while full fine-tuning was computationally prohibitive without providing significant performance gains.
Metrics
Standard metrics like Chamfer Distance (CD) often fail to capture topological defects in mechanical parts (e.g., a closed hole vs. an open hole). To address this, we utilize Exact Euler Characteristic Match (EECM) and Sphericity Discrepancy (SD). EECM directly reflects the number of holes and connected components, while SD measures the object’s deviation from a perfect sphere. Additionally, we employ the Discrete Mean Curvature Difference (DMCD) to calculate the average “mean curvature” across vertices. DMCD is useful for scenarios where surface curvature fidelity is crucial, such as comparing the smoothness of mechanical fillets. Collectively, these metrics are particularly helpful when scaling towards complex 3D object generation, where intricate differences in geometry and topology are critical.
@@ Appendix F Licenses for External Assets
The DeepCAD dataset (Wu et al., 2021) is available under the MIT License, as indicated in its official repository. The ABC dataset (Koch et al., 2019), which is a source for DeepCAD, is also distributed under the MIT License. The Fusion 360 Reconstruction dataset (Willis et al., 2021), originating from the Autodesk Fusion 360 Gallery, is referenced in academic literature as being available for research purposes under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.
The language models leveraged include: Qwen2.5-Coder (Yang et al., 2024), open-source versions of which are provided under the Apache License 2.0. The GPT-4.1 model (OpenAI, 2025) is a proprietary offering from OpenAI, and its usage is governed by OpenAI’s prevailing Terms of Use, Usage Policies, and Sharing & Publication Policy. Gemma models (Team et al., 2025), developed by Google, are released under a custom Gemma license.
Key software utilized includes Blender (Blender Online Community, 2025), which is distributed under the GNU General Public License (GPL) version 3 or later. The Python library trimesh (Dawson-Haggerty et al., ), used for geometric computations, is available under the MIT License.